OpenStudy (anonymous):

the acceleration of an equipped with a standard engine is: a = f(t) = 4 + .08t (0≤t≤12) t seconds after starting from rest at full throttle. The acceleration of an engine equipped with a turbo-charger is given by a = g(t) = 4 + 1.2t + 0.03t² (0≤t≤12). how much faster is the turbo charged model moving than the model with the standard engine at the end of a 10-sec test run at full throttle?

OpenStudy (anonymous):

x= double integral of a= (2t^2)+(.08/6)t^(3), standard engine x=(2t^2)+(1.2/6)t^(3)+(.03/12)t^(4), turbocharged engine after 10 seconds find which engine went further by letting t=10 in these equations so.. x=(400)+(80/6)=413.3, standard engine x=(400)+(200)+(300/12)=625, turbocharged engine

OpenStudy (anonymous):

let me know if this helped...

OpenStudy (anonymous):

okay I am going to look over it at 10 AM when I have free time. How did you find the first part though?

OpenStudy (anonymous):

if u integrate acceleration twice you can get the displacement...